### v1

# Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas

Abstract
The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic in curvature tensors U and V. Several classification results for gradient-shrinking Ricci solitons have been obtained under the assumption that the Bach tensor vanishes. We define a Bach-like tensor to be any other linear combination of U and V. We prove that, except along the Bach line, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton. Finally, we show that Bach-like tensors arise as Euler–Lagrange equations of a two-parameter family of quadratic curvature functionals and compute the corresponding first and second variation formulas.
1 Introduction and Main Results
A gradient Ricci soliton is a Riemannian manifold $(M^{n},g)$ together with a smooth potential function $f:M\to\mathbb{R}$ satisfying
$\mathrm{Rc}+\nabla^{2}f\;=\;\rho\,g,$ (1)
for $\rho$ a constant. The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience. Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities. Unless otherwise stated, our work will be in dimension $n=4$.
The classical Bach tensor was introduced by R. Bach [bach:1921] to study conformal relativity. It is defined in dimension 4 by:
$\mathrm{B}_{ij}=\nabla^{k}\nabla^{l}\mathrm{W}_{ijkl}+\frac{1}{2}\mathrm{R}^{kl}\mathrm{W}_{ikjl},$ (2)
where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $R_{ij}$ the Ricci curvature. It is well-known that $\mathrm{B}_{ij}$ is conformally invariant, symmetric, trace-free, divergence-free, and quadratic in the Riemann curvature tensor. As seen in [bergman:2004] there are exactly two independent symmetric 22-tensors that are divergence-free and quadratic in curvature in four dimensions. These are the following:
$\mathrm{U}_{ij}=2\mathrm{R}_{ipjq}\mathrm{R}^{pq}+\Delta\mathrm{R}_{ij}-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{1}{4}\mathrm{R}^{2}\,g_{ij},$ (3)
$\mathrm{V}_{ij}=-\nabla_{i}\nabla_{j}\mathrm{R}+\Delta\mathrm{R}\,g_{ij}+\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}\,g_{ij}.$ (4)
One can express the Bach tensor as a combination of $U$ and $V$. In dimension 4 we have:
$\mathrm{B}_{ij}=\frac{1}{2}\mathrm{U}_{ij}+\frac{1}{6}\mathrm{V}_{ij}.$ (5)
We define a Bach-like tensor to be of the form:
$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ (6)
for constants $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. If $\alpha/\beta=3$ then the Bach-like tensor is simply a rescaled Bach tensor. H.-D. Cao and Q. Chen showed in [caochen:2013] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid. Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity, but without the index-heavy computations typical of the Bach-flat case. On the one hand, our results here are somewhat stronger in that only two classifications occur; on the other, this suggests that the Bach tensor behaves more like a genuine analytic obstruction, requiring more delicate analysis that ties directly into the geometry of the conformal Cotton tensor $\mathrm{D}_{ijk}$, which will be discussed later.
Theorem 1.1 (Vanishing Basis-Tensors).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
  1. (i)
    $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton;
  2. (ii)
    $\mathrm{U}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
Theorem 1.2 (Bach-like Rigidity).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ such that $\beta\neq\alpha/3$. Then if
$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0,$
$M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathfrak{B}_{ij}=0$. Hence, within the class of complete gradient-shrinking Ricci solitons, $\mathfrak{B}_{ij}=0$ is equivalent to $M^{4}$ being Einstein or isometric to the Gaussian soliton.
We will approach the proof of Theorem 1.2 by first considering the cone
$\mathcal{C}=\{(\alpha,\beta)\,\,|\,\,\alpha\geq 0\,\,\mathrm{and}\,\,\beta>\alpha/3\}\cup\{(\alpha,\beta)\,\,|\,\,\alpha\leq 0\,\,\mathrm{and}\,\,\beta<\alpha/3\},$
and primarily using weighted-integral identities from the $\mathrm{V}_{ij}$ tensor. For the region outside the cone we will take a different approach using more general weighted-integral identities. This result covers the entire $(\alpha,\beta)$ plane aside from the Bach-line and the origin.
2 Integral Identities
To proceed, we rely on a number of integral identities implied by $\Delta R=0$. The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of U and V.
Lemma 2.1.
Let $U_{ij}$ and $V_{ij}$ be given by (3) and (4) respectively. Then
$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta R$
$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta R.$
So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta R=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
Lemma 2.2 (Integral Identities).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta\mathrm{R}=0$. Let $r$ be a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. Then:
  1. 1.)
    $\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S=0,$
  2. 2.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V=0,$
  3. 3.)
    $\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S,$
  4. 4.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V=0,$
  5. 5.)
    $\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=\frac{1}{2}\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V,$
  6. 6.)
    $\int_{\Omega_{r}}f\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V=0.$
Proof.
First, we have:
$0=\int_{\Omega_{r}}\Delta R\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla R,\nabla f\rangle\,\mathrm{d}S$
showing (1). Now we integrate by parts to get:
$\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V$ $=\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V$
$=r\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V.$
As $\Delta R=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $R\Delta R$:
$0=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
And by using the identity $\mathrm{R}=f-|\nabla f|^{2}$ from normalizing $f$ as in [hamilton:1988], we get (3). Now we have:
$0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{e}^{-f}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\mathrm{e}^{-r}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
showing (4). To show (5) we integrate $R\Delta R$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
$0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\langle\nabla\big(\mathrm{R}\,\mathrm{e}^{-f}\big),\nabla\mathrm{R}\rangle\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-R\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-2\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V.$
And lastly for (6), we have:
$\int_{\Omega_{r}}f\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\int_{\Omega_{r}}\langle f\nabla\,\mathrm{R},\nabla\mathrm{e}^{-f}\rangle\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-f}\,\mathrm{d}V$
$=-r\mathrm{e}^{-r}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\,\Delta\mathrm{R}\Big)\mathrm{e}^{-f}\,\mathrm{d}V$
$=0.$
3 The Vanishing V-tensor
The tensor $\mathrm{V}_{ij}$ has fewer and less complicated terms than $\mathrm{U}_{ij}$. We will obtain integral identities by assuming $\mathrm{V}_{ij}=0$, and use these to prove Theorem 1.1. We first consider the Hessian term:
Lemma 3.1.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta R=0$. Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
$\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\mathrm{e}^{-r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Proof.
We integrate:
$\int_{\Omega_{r}}\nabla^{2}$ $\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\langle\nabla\langle\nabla\mathrm{R},\nabla f\rangle,\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{2}\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\langle\nabla\langle\nabla\mathrm{R},\nabla f\rangle,\nabla\mathrm{e}^{-f}\rangle\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}\Delta(\mathrm{e}^{-f})\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}(|\nabla f|^{2}-\Delta f)\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}(f-2)\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Lemma 7.1 (1) and (3) shows that the integral of $(f-2)\langle\nabla\mathrm{R},\nabla f\rangle$ vanishes, giving us our desired result. ∎
It will be necessary to control the $|\nabla\mathrm{R}|^{2}$ integral as we allow $r\to\infty$. Work already exists for controlling the integral of $|\mathrm{Rc}|^{2}$, which we are able to translate to our problem.
Lemma 3.2.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then:
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V<\infty$
for any $\alpha>0$.
Proof.
By the Cauchy-Schwarz inequality we have:
$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}$
By H.-D. Cao and D. Zhou’s growth estimates for the potential function [caozhou:2010], we have:
$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2}$
[…]
$\int_{M}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V=\int_{K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V+\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V.$
Clearly the first integral over the compact set $K$ is finite. Now we have for the second integral:
$\begin{split}\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V&=\int_{M\backslash K}(|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2})|\nabla f|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\\ &\leq\int_{M\backslash K}|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.\end{split}$
By O. Munteanu and N. Sesum’s results in [munteanusesum:2013], we have:
$\int_{M}|\mathrm{Rc}|^{2}\mathrm{e}^{-\lambda f}\,\mathrm{d}V<\infty,$
for any $\lambda>0$, so after setting $\lambda=\frac{\alpha}{2}$ the result follows. ∎
Lemma 3.3.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then if $\Omega_{r}=\{f\leq r\}$ for regular values $r>0$ of $f$ and $\alpha>0$:
$\lim_{r\rightarrow\infty}\Big(\mathrm{e}^{-\alpha r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)=0.$
Proof.
We first apply Lemma 3.2:
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty$
Since $f\leq r$ on $\Omega_{r}$, we also have $\mathrm{e}^{-\alpha r/2}\leq\mathrm{e}^{-\alpha f/2}$:
$\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V=\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha r/2}\,\mathrm{d}V\leq\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.$
Now let $r\rightarrow\infty$, giving us the desired result:
$\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)\leq\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\Big)=0.$
4 Proof of Theorem 1.1(i)
Proof.
Let $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$ where $r$ is a regular value of $f$. By Lemma 2.1, $\mathrm{V}\equiv 0$ implies $\Delta\mathrm{R}=0$. By Lemma 2.2(5) and (3) we have:
$\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\Delta_{f}\mathrm{R}\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-r}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
And using Lemma 3.1, we compute the integral of $V(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, and we note that the second term involving $\Delta R$ again vanishes:
$\int_{\Omega_{r}}\mathrm{V}($ $\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\Big(\nabla^{2}\mathrm{R}(\nabla f,\nabla f)-\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)+\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\frac{\mathrm{e}^{-r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Letting $r\rightarrow\infty$, by Lemma 3.3 the first integral vanishes. We obtain:
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\int_{M}\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$ (7)
Assuming $\mathrm{V}_{ij}=0$, for all $p\in M$ we must have either $\mathrm{R}=0$ or $|\nabla f|=0$. By the work of S. Pigola, M. Rimoldi, and A. Setti [pigolarimoldi:2011], either $R>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $R=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$.
For the converse, suppose $(M^{4},g,f)$ is Einstein. Then as the scalar curvature is constant, we have:
$\mathrm{V}_{ij}$ $=\frac{1}{4}\mathrm{R}^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}$
$=-\mathrm{R}(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij})$
$=0.$
If the manifold is Gaussian, then $\mathrm{R}=0$ and again $\mathrm{V}=0$. ∎
5 Weighted Bach Integral
In [caochen:2013] H.-D. Cao and Q. Chen consider how the Bach tensor links to the Cotton tensor of a conformal metric, called $\mathrm{D}_{ijk}$. The norm of $\mathrm{D}_{ijk}$ contains rich geometric information, and they were able to use it to prove their classification results. We will link Bach-like tensors to $\mathrm{D}_{ijk}$ as well. Recall that the Cotton tensor is given by:
$\mathrm{C}_{ijk}=-\frac{n-2}{n-3}\nabla_{l}\mathrm{W}_{ijkl},$
where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices. That is,
$\mathrm{C}_{ijk}=-\mathrm{C}_{jik}\quad\mathrm{and}\quad g^{ij}\,\mathrm{C}_{ijk}=g^{ik}\,\mathrm{C}_{ijk}=0.$
The Cotton tensor for the conformal metric $\widehat{g}_{ij}=\mathrm{e}^{\frac{1}{2}f}\,g_{ij}$ is given by:
$\mathrm{D}_{ijk}=\mathrm{C}_{ijk}+\mathrm{W}_{ijkl}\,\nabla_{l}f.$
On a complete gradient-shrinking Ricci soliton, $\mathrm{D}_{ijk}$ can be expressed as:
$D_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
$\quad-\frac{1}{(n-1)(n-2)}\mathrm{R}\,\big(g_{jk}\,\nabla_{i}f-g_{ik}\,\nabla_{j}f\big).$
H.-D. Cao and Q. Chen [caochen:2013] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
$\mathrm{B}_{ij}=-\frac{1}{n-2}\Big(\nabla_{k}\mathrm{D}_{ikj}+\frac{n-3}{n-2}\mathrm{C}_{jli}\nabla_{l}f\Big).$
We can now link together the integral of the Bach tensor contracted with the gradient of the potential function to the norm of $\mathrm{D}_{ijk}$.
Lemma 5.1.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
$\int_{M}\mathrm{B}(\nabla f,\nabla f)e^{-f}\,\mathrm{d}V=-\frac{1}{2}\int_{M}|\mathrm{D}_{ijk}|^{2}e^{-f}\,\mathrm{d}V.$
Proof.
Let $r$ be a regular value of $f$ and $\Omega_{r}=\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
$(n-2)\int_{\Omega_{r}}$ $\mathrm{B}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\nabla_{k}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-f}\,\mathrm{d}V-\frac{n-3}{n-2}\int_{\Omega_{r}}\mathrm{C}_{jki}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\nabla_{k}f\cdot|\nabla f|^{-1}\,\mathrm{e}^{-f}\,\mathrm{d}S$
$\quad+\int_{\Omega_{r}}\mathrm{D}_{ikj}(\mathrm{R}_{ik}\nabla_{j}f+\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$\quad-\int_{M}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{D}_{ikj}(\mathrm{R}_{ik}\nabla_{j}f+\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{1}{2}\int_{\Omega_{r}}\mathrm{D}_{ijk}(\mathrm{R}_{ik}\nabla_{j}f-\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{(n-2)}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
6 Proof of Theorem 1.2
First observe that for any $(\alpha,\beta)\in\mathcal{C}$, and a Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ we can write:
$\mathfrak{B}_{ij}=2\alpha\,\mathrm{B}_{ij}+\Big(\beta-\frac{\alpha}{3}\Big)\,\mathrm{V}_{ij}.$
Assuming $\mathfrak{B}_{ij}=0$ and taking a trace we have:
$g^{ij}\,\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0\implies\Delta\mathrm{R}=0,$
by Lemma 2.1. Then we have:
$\int_{M}\mathfrak{B}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\alpha\int_{M}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{4}\Big(\beta-\frac{\alpha}{3}\Big)\int_{M}\mathrm{R}^{2}\,|\nabla f|^{2}\,\mathrm{e}^{-f}\,\,\mathrm{d}V=0.$
Since $(\alpha,\beta)\in\mathcal{C}$, we have two cases. If $\alpha>0$ and $\beta>\alpha/3$, then we must have $|\mathrm{D}_{ijk}|^{2}(p)=0$ and $\mathrm{R}^{2}(p)|\nabla f(p)|^{2}=0$ for all $p\in M$. Then we can conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1. The argument for when $\alpha<0$ and $\beta<\alpha/3$ is analogous.
7 The Vanishing U-tensor
In order to extend the $(\alpha,\beta)$-cone in Theorem 1.2, we need to consider the implications of $\mathrm{U}_{ij}=0$. We attempt the same style of argument as we did with $\mathrm{V}_{ij}=0$. Define a weighted volume element of the form
$\mathrm{d}V_{c}(\mathbf{x})=\mathrm{e}^{-cf(\textbf{x})}\,\mathrm{d}V\quad\mathrm{where}\quad c>0.$
Lemma 7.1.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Let $r$ be a regular value of $f$ and define the sub-level set $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. If $c>0$ is a constant and $\Delta\mathrm{R}=0$, then we have the following:
  1. (1.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=0$,
  2. (2.)
    $\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$,
  3. (3.)
    $\int_{\Omega_{r}}f\,\langle\nabla R,\nabla f\rangle\,\mathrm{d}V_{c}=0$.
Proof.
First we have:
$0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S+c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V$
$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\,\mathrm{d}V.$
Next, we have:
$0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-c\,\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\Big)\mathrm{e}^{-cf}\,\mathrm{d}V.$
Hence,
$\int_{\Omega_{r}}\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
where we used 2.2(3). And the last critical identity is:
$\int_{\Omega_{r}}f$ $\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\frac{1}{c}\int_{\Omega_{r}}\langle f\,\nabla\mathrm{R},\nabla(\mathrm{e}^{-cf})\rangle\,\mathrm{d}V$
$=-\frac{1}{c}\Big[\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\mathrm{div}(f\nabla\mathrm{R})\mathrm{e}^{-cf}\,\mathrm{d}V\Big]$
$=-\frac{1}{c}\Big[r\,\mathrm{e}^{-cr}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-cf}\,\mathrm{d}V\Big]$
$=\frac{1}{c}\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\Delta\mathrm{R}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V$
$=0.$
We first state a quick consequence of Bochner’s formula, and then we will proceed with the main integral identity.
Lemma 7.2.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Then if $\Delta\mathrm{R}=0$:
$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle$
Proof.
By Bochner’s formula, we have:
$\frac{1}{2}\Delta|\nabla f|^{2}$ $=\langle\nabla(\Delta f),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\mathrm{Rc}(\nabla f,\nabla f)$
$\frac{1}{2}\Delta(f-\mathrm{R})$ $=\langle\nabla(2-\mathrm{R}),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$
$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\frac{1}{2}g-\mathrm{Rc}|^{2}$
$1-\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+1-\mathrm{R}+|\mathrm{Rc}|^{2}$
$\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\mathrm{Rc}|^{2}.$
From here our claim clearly follows. ∎
Lemma 7.3.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Moreover, this can be rewritten as
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
Proof.
One can easily check that on a soliton we have
$U_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
as well as the soliton identities
  1. (a)
    $\mathrm{\mathrm{Rc}}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{\mathrm{Rc}}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
  2. (b)
    $\Delta f=2-\mathrm{R}$,
  3. (c)
    $\nabla^{2}f(\nabla f)=\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}$.
The second term in our expression for $U$ is simplified by looking at the following divergence:
$\nabla_{k}(\mathrm{R}_{ij}$ $\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf})$
$=\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\mathrm{e}^{-cf}+2\,\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}\nabla_{k}f(\nabla_{k}f)\mathrm{e}^{-cf}$
$\quad+\Delta f\,\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-cf}-c|\nabla f|^{2}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf}.$
Now by the divergence theorem we have
$\int_{\Omega_{r}}$ $\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{div}(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
$\quad-\int_{\Omega_{r}}\langle\nabla\mathrm{R},\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\frac{1}{2}\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S+\frac{1}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
So the weighted integral of $\mathrm{U}_{ij}$ becomes
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Now we can transform the integral using our Bochner formula from Lemma 7.2 as well as
$\int_{\Omega_{r}}$ $\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}$
$=\frac{1}{4}\int_{\Omega_{r}}\Big(\mathrm{R}^{2}|\nabla f|^{2}-\mathrm{R}|\nabla f|^{2}+\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{d}V_{c}$
$=\frac{1}{4}\int_{\Omega_{r}}\Big(\mathrm{R}^{2}|\nabla f|^{2}-\mathrm{R}|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
The issue is that $\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}$ can change sign over $M$, so
$\int_{M}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}=0$
does not yield any obvious rigidity. The $\mathrm{R}^{2}|\nabla f|^{2}$ term reminds us of our work on the $\mathrm{V}$-tensor, but we must first reformulate our integral identity for $\mathrm{V}_{ij}$ for the altered-weighted volume element.
Lemma 7.4.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. Then if $\Delta\mathrm{R}=0$ and $c>0$, we have
$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
and
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}.$
Proof.
We will once again integrate the terms of $\mathrm{V}(\nabla f,\nabla f)$ separately. First, we compute the following divergence:
$\nabla_{i}$ $(\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf})$
$=\nabla_{i}\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-cf}+\Delta f\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}$
$\quad+\nabla^{2}f(\nabla f,\nabla\mathrm{R})\mathrm{e}^{-cf}-c|\nabla f|^{2}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}.$
Hence, the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
$\int_{\Omega_{r}}$ $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\nabla_{i}(\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{i}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\int_{\Omega_{r}}(\mathrm{R}-2)\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\int_{\Omega_{r}}\Big(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle-\frac{1}{2}|\nabla\mathrm{R}|^{2}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V+c\int_{\Omega_{r}}(f-\mathrm{R})\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad+(c-1)\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\Big(\frac{c-1}{c}+\frac{1}{2}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V.$
Therefore, we have:
$\int_{\Omega_{r}}$ $\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\Big(\frac{c-1}{c}+\frac{1}{2}-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
Now taking the limit $r\rightarrow\infty$:
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V.$
8 Proof of Theorem 1.1(ii)
Assume $\mathrm{U}_{ij}\equiv 0$. By Lemma 2.1, $\mathrm{tr}_{g}(\mathrm{U})=-\Delta\mathrm{R}$, hence $\Delta R\equiv 0$. Fix a regular value $r\in(0,1)$ of $f$ and write $\Omega_{r}=\{x\in M:f(x)\leq r\}$. For any $c>0$, Lemma 7.3 gives
$0=\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\!\Big((\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}\!|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
Multiplying by $4c$ and rearranging yields the basic identity
$c\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\big(\mathrm{e}^{-cr}-\mathrm{e}^{-cf}\big)\,\,\mathrm{d}V.$
On the fixed bounded set $\Omega_{r}$ we have $0\leq f\leq r<1$, so all integrands are smooth and dominated by an $L^{1}$ function independent of $c$. Hence we may differentiate under the integral sign. Differentiating once gives
$\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V-c\int_{\Omega_{r}}f\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(f\,\mathrm{e}^{-cf}-r\,\mathrm{e}^{-cr}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
Differentiating again yields
$\int_{\Omega_{r}}\big(cf^{2}-2f\big)(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{2}\,\mathrm{e}^{-cr}-f^{2}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
An easy induction shows that for each integer $n\geq 0$,
$\int_{\Omega_{r}}\big(nf^{\,n-1}-cf^{\,n}\big)\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{\,n}\,\mathrm{e}^{-cr}-f^{\,n}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$ (8)
Setting $c=0$ in (8) and summing over $n\geq 0$, we use the geometric series sum formulas:
$\sum_{n=0}^{\infty}nf^{\,n-1}=\frac{1}{(1-f)^{2}},\qquad\sum_{n=0}^{\infty}f^{\,n}=\frac{1}{1-f},\qquad\sum_{n=0}^{\infty}r^{\,n}=\frac{1}{1-r}.$
This leads to the identity
$\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.$ (9)
On $\Omega_{r}$ we have $0\leq\mathrm{R}\leq f<1$, hence $\mathrm{R}^{2}-\mathrm{R}\leq 0$ and $f-r\leq 0$, while (1f)>0(1-f)>0 and (1r)>0(1-r)>0. Therefore both terms in the integrand of (9) are non-positive. Since the integral vanishes, both terms vanish identically on $\Omega_{r}$:
[…]
In particular $\nabla\mathrm{R}\equiv 0$ on $\Omega_{r}$. Together with $\Delta\mathrm{R}\equiv 0$, unique continuation implies $\mathrm{R}$ is constant on all of $M$. If $\mathrm{R}=0$ at some point then $\mathrm{R}\equiv 0$ everywhere and the soliton is isometric to the Gaussian soliton on $\mathbb{R}^{4}$. Otherwise $\mathrm{R}>0$ and the soliton is Einstein.
9 Proof of Theorem 1.2 Outside the Cone
Assume $3\beta\neq\alpha$, then $\mathrm{tr}_{g}\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0$, hence $\Delta\mathrm{R}\equiv 0$. For any $c>0$ and a fixed regular value $r\in(0,1)$ for $f$, let $\Omega_{r}=\{f\leq r\}.$ We express the Bach-like tensor as
$\mathfrak{B}_{ij}=6\beta\,\mathrm{B}_{ij}+(\alpha-3\beta)\,\mathrm{U}_{ij}$
One can easily check that for the volume element $\mathrm{d}V_{c}$, the integration of the Bach tensor yields:
$\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=-\frac{c}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}=\mathrm{F}(c).$
So differentiating past the integral symbol yields
$\frac{d^{n}}{dc^{n}}\Big|_{c=0}\mathrm{F}(c)=-\frac{1}{2}\int_{\Omega_{r}}nf^{n-1}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V.$
We will need to multiply both sides of the Bach-like equation by $c$ to proceed.
$\frac{d^{n}}{dc^{n}}\Big(c\,\mathrm{F}(c)\Big)=n\,\mathrm{F}^{(n-1)}(c)+c\,\mathrm{F}^{(n)}(c),$
hence
$\frac{d^{n}}{dc^{n}}\Big|_{c=0}=n\,\mathrm{F}^{(n-1)}(0)=-\frac{1}{2}\int_{\Omega_{r}}n(n-1)\,f^{n-2}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V$
Summing over $n$ on the right-hand side, we obtain
$\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V.$ (10)
Now assuming $\mathfrak{B}_{ij}\equiv 0$, and multiplying both sides by $c$:
$0=6\beta\,c\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c},$
hence
$0=\frac{-6\beta c^{2}}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
Now differentiating with respect to $c$, evaluating each derivative at $c=0$ and summing over $n$ we finally obtain
$\boxed{-6\beta\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V+\frac{1}{4}(\alpha-3\beta)\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.}$
Now as $0<r<1$, both terms in the second integrand are negative as in (9). If $\beta>0$ and $\alpha-3\beta>0$, or if $\beta<0$ and $\alpha-3\beta<0$, then
$|\mathrm{D}_{ijk}|^{2}\equiv 0,\quad\mathrm{R}^{2}-\mathrm{R}\equiv 0,\quad\mathrm{and}\quad|\nabla\mathrm{R}|^{2}\equiv 0\quad\mathrm{on}\,\,\Omega_{r}.$
Hence, $M$ must be Einstein or isometric to the Gaussian shrinking soliton. These ranges for $\alpha$ and $\beta$ give us the desired result for the region outside the cone $\mathcal{C}$, hence for all $(\alpha,\beta)$ aside from $(0,0)$ and the Bach-line.
10 Variational Origins of U and V
We now connect the tensors U and V with the variational structure of the Riemannian metric. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. Throughout, we work in the standard transverse–traceless (TT) gauge, imposing for a variation $h$
$\mathrm{(TT)\,\,gauge:}\quad\mathrm{Div}_{g}(h)=0\quad\mathrm{and}\quad\mathrm{tr}_{g}(h)=0.$
Lemma 10.1.
In dimension $4$, define
$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|W_{g}|^{2}\,dV_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}R_{g}^{2}\,dV_{g}\,.$
Then the $L^{2}$–gradient satisfies
$\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}\;=\;\alpha\,\mathrm{U}\;+\;\beta\,\mathrm{V}.$
In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is equivalent to the vanishing of the Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
Proof.
In dimension four, one has
$\delta\!\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{B},\delta g\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta\!\int\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{V},\delta g\rangle\,\mathrm{d}V_{g}.$
Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.$
Hence $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$. ∎
Corollary 10.2.
$\mathrm{V}\;=\;-\tfrac{1}{2}\,\mathrm{grad}\!\left(\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}\right),\qquad\mathrm{U}\;=\;-\,\mathrm{grad}\!\left(\int_{M}\big(|\mathrm{W}|^{2}-\tfrac{1}{6}\mathrm{R}^{2}\big)\,\,\mathrm{d}V_{g}\right).$
Proposition 10.3 (Second variation of $\int R^{2}$ in TT gauge).
Let $(M^{4},g)$ be closed. For a variation $g_{t}=g+t\,h+o(t)$ with
$\mathrm{div}h=0,\qquad\mathrm{tr}_{g}h=0\quad\text{\emph{(TT gauge)}},$
the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}\;-\;2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}\;+\;2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}\;+\;2\langle\mathrm{Rc},h\rangle^{2}\;+\;\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,dV\,.\;$
Proof.
Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Then by differentiating once more, and using $\mathrm{div}h=0$, $\mathrm{tr}_{g}h=0$, the Hessian is
$\delta^{2}F[h,h]=-2\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
We have the following standard TT-gauge variations:
$\delta\mathrm{R}$ $=-\langle\mathrm{Rc},h\rangle,$
$\delta(\nabla_{i}\nabla_{j}\mathrm{R})$ $=\nabla_{i}\nabla_{j}(\delta\mathrm{R})-\tfrac{1}{2}(\nabla_{i}h_{j}{}^{k}+\nabla_{j}h_{i}{}^{k}-\nabla^{k}h_{ij})\,\nabla_{k}\mathrm{R},$
$\delta(\Delta\mathrm{R})$ $=\Delta(\delta\mathrm{R})-\langle h,\nabla^{2}\mathrm{R}\rangle,$
$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij}.$
Using these, the Hessian can be written as
$\delta^{2}F[h,h]=\int_{M}\Big(-\mathrm{R}\,\langle\Delta_{L}h,h\rangle+2\langle\mathrm{Rc},h\rangle^{2}+(\frac{1}{2}\mathrm{R}^{2}-2\Delta\mathrm{R})|h|^{2}\Big)\,\,\mathrm{d}V_{g}.$
We expand the Lichnerowicz laplacian as $\Delta_{L}h=\nabla^{*}\nabla h+2\mathrm{Rm}\!*h-\mathrm{Rc}\circ h-h\circ\mathrm{Rc}$. Integration by parts yields the result. ∎
Remark 10.4.
If $(M^{4},g)$ is Einstein, the main contributions to the second variation come from the derivative term $\mathrm{R}\,|\nabla h|^{2}$ and the curvature terms involving $\mathrm{R}_{ikjl}\,h_{ij}\,h_{kl}$.
Proposition 10.5 (Second variation at Einstein backgrounds).
On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and scalar curvature $R=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in the TT-gauge is
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle dV.$
Equivalently, for each TT eigenmode $\Delta_{L}h_{k}=\mu_{k}h_{k}$,
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\sum_{k}p_{\alpha,\beta}(\mu_{k})\,\|h_{k}\|_{L^{2}}^{2},\quad p_{\alpha,\beta}(\mu)=\alpha\mu^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\mu+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta).$
Proof.
In TT gauge at an Einstein metric, the standard linearizations give
$\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\qquad\mathrm{V}^{\prime}(h)=\frac{1}{2}\mathrm{R}\,\Delta_{L}h.$
Since
$\delta^{2}\!\int_{M}|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta^{2}\!\int_{M}\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g},$
the claimed polynomial follows by expanding
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
Corollary 10.6 (Bach line and factorization).
On the Bach line $\beta=\alpha/3$, one has
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g},$
i.e. the Hessian factorizes into two shifted Lichnerowicz factors.
Corollary 10.7 (Local rigidity by spectral positivity).
Let $(M^{4},g_{E})$ be Einstein with TT–spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$. If $p_{\alpha,\beta}(\mu)>0$ for every TT–eigenvalue $\mu$ (equivalently, $\inf_{\mu\geq\mu_{0}}p_{\alpha,\beta}(\mu)>0$), then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling. In particular, there exists a neighborhood of $g_{E}$ in which the equation $\alpha U+\beta V=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
Remark 10.8 (Flat/Gaussian background).
If $R=0=W$ (Euclidean space, or a flat compact quotient), then $B^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $V^{\prime}(h)\equiv 0$ on TT. Hence
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,dV=\alpha\int_{M}|\Delta h|^{2}\,dV\quad(\text{TT}),$
which is strictly positive for $\alpha>0$ on compactly supported (or suitably decaying/weighted) variations.