### v3

# arXiv:2511.05774v3[math.DG] 02 Mar 2026

Abstract
The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic-in-curvature tensors $\mathrm{U}$ and $\mathrm{V}$. Several classification results for gradient-shrinking Ricci solitons have been obtained under the assumption that the Bach tensor vanishes. We define a Bach-like tensor to be any other linear combination of $\mathrm{U}$ and $\mathrm{V}$. We prove that within a certain cone of parameters $\mathcal{C}$, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton, extending the results of Cao–Chen [4]. The special case where $\mathrm{U}=0$ forces $f_{\min}\in\{0,1,2\}$, with rigidity holding when $f_{\min}=0,2$. The remaining case $f_{\min}=1$ is the central open problem, with $\mathbb{R}^{2}\times\mathbb{S}^{2}(\sqrt{2})$ as the conjectured exceptional geometry. Finally, we show that Bach-like tensors arise as Euler–Lagrange equations of a two-parameter family of quadratic curvature functionals and compute the corresponding first and second variation formulas.
1 Introduction and Main Results
A gradient Ricci soliton is a Riemannian manifold $(M^{n},g)$ together with a smooth potential function $f:M\to\mathbb{R}$ satisfying
$\mathrm{Rc}+\nabla^{2}f\;=\;\rho\,g,$ (1)
for $\rho$ a constant. The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience. We normalize $f$ throughout so that
$\mathrm{R}+|\nabla f|^{2}=f,$ (2)
following Hamilton [6]. Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities. Unless otherwise stated, our work will be in dimension $n=4$.
The classical Bach tensor was introduced by R. Bach [2] to study conformal relativity. It is defined in dimension 4 by:
$\mathrm{B}_{ij}=\nabla^{k}\nabla^{l}\mathrm{W}_{ijkl}+\frac{1}{2}\mathrm{R}^{kl}\mathrm{W}_{ikjl},$ (3)
where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $\mathrm{R}_{ij}$ the Ricci curvature. It is well-known that $\mathrm{B}_{ij}$ is conformally invariant, symmetric, trace-free, divergence-free, and quadratic in the Riemann curvature tensor. As seen in [1] there are exactly two independent symmetric $2$-tensors that are divergence-free and quadratic in curvature in four dimensions:
$\mathrm{U}_{ij}=2\mathrm{R}_{ipjq}\mathrm{R}^{pq}+\Delta\mathrm{R}_{ij}-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{1}{4}\mathrm{R}^{2}\,g_{ij},$ (4)
$\mathrm{V}_{ij}=-\nabla_{i}\nabla_{j}\mathrm{R}+\Delta\mathrm{R}\,g_{ij}+\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}\,g_{ij}.$ (5)
One can express the Bach tensor as a combination of $\mathrm{U}$ and $\mathrm{V}$. In dimension 4:
$\mathrm{B}_{ij}=\frac{1}{2}\mathrm{U}_{ij}+\frac{1}{6}\mathrm{V}_{ij}.$ (6)
We define a Bach-like tensor to be of the form:
$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ (7)
for constants $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. If $\alpha/\beta=3$ then the Bach-like tensor is simply a rescaled Bach tensor. H.-D. Cao and Q. Chen showed in [4] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid. Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity within a certain parameter cone, but without the index-heavy computations typical of the Bach-flat case. The principal results of this paper are as follows.
Theorem 1.1 (Vanishing V-tensor).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Then $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. In particular, $f_{\min}\in\{0,2\}$ correspondingly.
Theorem 1.2 (Vanishing U-tensor).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\mathrm{U}_{ij}\equiv 0$ then $f_{\min}\in\{0,1,2\}$. If $f_{\min}=0$ then $M^{4}$ is isometric to the Gaussian shrinking soliton, and if $f_{\min}=2$ then $M^{4}$ is Einstein. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathrm{U}_{ij}\equiv 0$.
The case $f_{\min}=1$ in Theorem 1.2 is not resolved by the methods of this paper. At the minimum point $p$ of $f$, the tensor $\mathrm{U}$ vanishes independently of any global assumption when $\mathrm{R}(p)=1$, so the local analysis yields no information. By contrast, $\mathrm{V}$ cannot vanish at $p$ when $f_{\min}=1$. We state the following as a conjecture.
Conjecture 1.3.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\mathrm{U}\equiv 0$ and $f_{\min}=1$, then $M^{4}$ is isometric to $\mathbb{R}^{2}\times S^{2}(\sqrt{2})$.
The candidate space $\mathbb{R}^{2}\times S^{2}(\sqrt{2})$ is verified to satisfy the hypotheses. For the shrinking soliton normalization $\rho=\tfrac{1}{2}$, the soliton equation in the $S^{2}$ directions forces radius $r=\sqrt{2}$, giving $\mathrm{R}=1$ and $f_{\min}=1$ via (2). Indexing by $0,1$ in the $\mathbb{R}^{2}$-directions and $a,b$ on $S^{2}(\sqrt{2})$, a direct computation gives $\mathrm{U}_{00}=\mathrm{U}_{0a}=\mathrm{U}_{ab}=0$, so $\mathrm{U}\equiv 0$. Moreover, $\mathrm{V}_{00}=-\tfrac{1}{4}$ and $\mathrm{V}_{ab}=\tfrac{1}{4}g_{ab}$, so $\mathrm{V}\not\equiv 0$, confirming the space is neither Einstein nor Gaussian. This computation was suggested to us by H.-D. Cao.
Theorem 1.4 (Bach-like Rigidity).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Assume $(\alpha,\beta)\in\mathcal{C}$ where
$\mathcal{C}=\{(\alpha,\beta)\,\,|\,\,\alpha\geq 0\,\,\mathrm{and}\,\,\beta>\alpha/3\}\cup\{(\alpha,\beta)\,\,|\,\,\alpha\leq 0\,\,\mathrm{and}\,\,\beta<\alpha/3\},$
and assume $3\beta\neq\alpha$. If $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$, then $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathfrak{B}_{ij}=0$. If $3\beta=\alpha$, the rigidity results of Cao–Chen [4] apply.
We approach the proof of Theorem 1.4 by primarily using weighted integral identities from the $\mathrm{V}_{ij}$ tensor inside the cone $\mathcal{C}$. This result covers the entire $(\alpha,\beta)$ plane aside from the Bach line, the $\beta=0$ axis, and the origin; a complete treatment of the remaining cases is contingent on Conjecture 1.3 and is deferred to future work.
2 Integral Identities
To proceed, we rely on a number of integral identities implied by $\Delta\mathrm{R}=0$. The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of $\mathrm{U}$ and $\mathrm{V}$.
Lemma 2.1.
Let $\mathrm{U}_{ij}$ and $\mathrm{V}_{ij}$ be given by (4) and (5) respectively. Then
$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta\mathrm{R},$
$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta\mathrm{R}.$
So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta\mathrm{R}=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
Lemma 2.2 (Integral Identities).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2), and assume that $\Delta\mathrm{R}=0$. Let $r$ be a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. Then:
  1. 1.)
    $\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S=0,$
  2. 2.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V=0,$
  3. 3.)
    $\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S,$
  4. 4.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V=0,$
  5. 5.)
    $\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=\frac{1}{2}\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V,$
  6. 6.)
    $\int_{\Omega_{r}}f\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V=0.$
Proof.
First, we have:
$0=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S$
showing (1). Now we integrate by parts to get:
$\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V$ $=\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V$
$=r\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V.$
As $\Delta\mathrm{R}=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $\mathrm{R}\Delta\mathrm{R}$:
$0=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
And by using the normalization (2), we get (3). Now we have:
$0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{e}^{-f}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\mathrm{e}^{-r}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
showing (4). To show (5) we integrate $\mathrm{R}\Delta\mathrm{R}$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
$0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\langle\nabla\big(\mathrm{R}\,\mathrm{e}^{-f}\big),\nabla\mathrm{R}\rangle\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-2\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V.$
And lastly for (6), we have:
$\int_{\Omega_{r}}f\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\int_{\Omega_{r}}\langle f\nabla\,\mathrm{R},\nabla\mathrm{e}^{-f}\rangle\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-f}\,\mathrm{d}V$
$=-r\mathrm{e}^{-r}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\,\Delta\mathrm{R}\Big)\mathrm{e}^{-f}\,\mathrm{d}V$
$=0.$
3 The Vanishing V-tensor
The tensor $\mathrm{V}_{ij}$ has fewer and less complicated terms than $\mathrm{U}_{ij}$. We will obtain integral identities by assuming $\mathrm{V}_{ij}=0$, and use these to prove Theorem 1.1. We first consider the Hessian term.
Lemma 3.1.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2), and assume that $\Delta\mathrm{R}=0$. Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
$\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\mathrm{e}^{-r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Proof.
We integrate:
$\int_{\Omega_{r}}\nabla^{2}$ $\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\langle\nabla\langle\nabla\mathrm{R},\nabla f\rangle,\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{2}\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\langle\nabla\langle\nabla\mathrm{R},\nabla f\rangle,\nabla\mathrm{e}^{-f}\rangle\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}\Delta(\mathrm{e}^{-f})\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}(|\nabla f|^{2}-\Delta f)\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}(f-2)\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Lemma 7.1(1) and (3) shows that the integral of $(f-2)\langle\nabla\mathrm{R},\nabla f\rangle$ vanishes, giving us our desired result. ∎
It will be necessary to control the $|\nabla\mathrm{R}|^{2}$ integral as we allow $r\to\infty$. Work already exists for controlling the integral of $|\mathrm{Rc}|^{2}$, which we are able to translate to our problem.
Lemma 3.2.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then:
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V<\infty$
for any $\alpha>0$.
Proof.
By the Cauchy-Schwarz inequality we have:
$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}.$
By H.-D. Cao and D. Zhou’s growth estimates for the potential function [5], we have:
$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2},$
[…]
$\int_{M}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V=\int_{K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V+\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V.$
The first integral over the compact set $K$ is finite. We show that the second integral is also finite:
$\begin{split}\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V&=\int_{M\backslash K}(|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2})|\nabla f|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\\ &\leq\int_{M\backslash K}|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.\end{split}$
By O. Munteanu and N. Sesum’s results in [7], we have:
$\int_{M}|\mathrm{Rc}|^{2}\mathrm{e}^{-\lambda f}\,\mathrm{d}V<\infty,$
for any $\lambda>0$, so after setting $\lambda=\frac{\alpha}{2}$ the result follows. ∎
Lemma 3.3.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then if $\Omega_{r}=\{f\leq r\}$ for regular values $r>0$ of $f$ and $\alpha>0$:
$\lim_{r\rightarrow\infty}\Big(\mathrm{e}^{-\alpha r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)=0.$
Proof.
We first apply Lemma 3.2:
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty.$
Since $f\leq r$ on $\Omega_{r}$, we also have $\mathrm{e}^{-\alpha r/2}\leq\mathrm{e}^{-\alpha f/2}$:
$\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V=\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha r/2}\,\mathrm{d}V\leq\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.$
Now let $r\rightarrow\infty$, giving us the desired result:
$\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)\leq\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\Big)=0.$
4 Proof of Theorem 1.1
Proof.
Let $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$ where $r$ is a regular value of $f$. By Lemma 2.1, $\mathrm{V}\equiv 0$ implies $\Delta\mathrm{R}=0$. By Lemma 2.2(5) and (3) we have:
$\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\Delta_{f}\mathrm{R}\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-r}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
And using Lemma 3.1, we compute the integral of $\mathrm{V}(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, noting that the term involving $\Delta\mathrm{R}$ vanishes:
$\int_{\Omega_{r}}\mathrm{V}($ $\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\Big(\nabla^{2}\mathrm{R}(\nabla f,\nabla f)-\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)+\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\frac{\mathrm{e}^{-r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Letting $r\rightarrow\infty$, by Lemma 3.3 the first integral vanishes. We obtain:
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\int_{M}\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$ (8)
Assuming $\mathrm{V}_{ij}=0$, we must have either $\mathrm{R}=0$ or $|\nabla f|=0$ pointwise. By the work of S. Pigola, M. Rimoldi, and A. Setti [8], either $\mathrm{R}>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $\mathrm{R}=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$. The values $f_{\min}\in\{0,2\}$ follow from (2): if Einstein then $\nabla f\equiv 0$ and $\mathrm{Rc}=\tfrac{1}{2}g$, giving $\mathrm{R}=2$ and $f_{\min}=2$; if Gaussian then $\mathrm{R}=0$ and $f_{\min}=0$.
For the converse, suppose $(M^{4},g,f)$ is Einstein. Then as the scalar curvature is constant, we have:
$\mathrm{V}_{ij}$ $=\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}g_{ij}=-\mathrm{R}\Big(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij}\Big)=0.$
If the manifold is Gaussian, then $\mathrm{R}=0$ and again $\mathrm{V}=0$. ∎
5 Weighted Bach Integral
In [4] H.-D. Cao and Q. Chen consider how the Bach tensor links to the Cotton tensor of a conformal metric, called $\mathrm{D}_{ijk}$. The norm of $\mathrm{D}_{ijk}$ contains rich geometric information, and they were able to use it to prove their classification results. We will link Bach-like tensors to $\mathrm{D}_{ijk}$ as well. Recall that the Cotton tensor is given by:
$\mathrm{C}_{ijk}=-\frac{n-2}{n-3}\nabla_{l}\mathrm{W}_{ijkl},$
where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices:
$\mathrm{C}_{ijk}=-\mathrm{C}_{jik}\quad\mathrm{and}\quad g^{ij}\,\mathrm{C}_{ijk}=g^{ik}\,\mathrm{C}_{ijk}=0.$
The Cotton tensor for the conformal metric $\widehat{g}_{ij}=\mathrm{e}^{\frac{1}{2}f}\,g_{ij}$ is given by:
$\mathrm{D}_{ijk}=\mathrm{C}_{ijk}+\mathrm{W}_{ijkl}\,\nabla_{l}f.$
On a complete gradient-shrinking Ricci soliton, $\mathrm{D}_{ijk}$ can be expressed as:
$\mathrm{D}_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
$\quad-\frac{1}{(n-1)(n-2)}\mathrm{R}\,\big(g_{jk}\,\nabla_{i}f-g_{ik}\,\nabla_{j}f\big).$
H.-D. Cao and Q. Chen [4] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
$\mathrm{B}_{ij}=-\frac{1}{n-2}\Big(\nabla_{k}\mathrm{D}_{ikj}+\frac{n-3}{n-2}\mathrm{C}_{jli}\nabla_{l}f\Big).$
Lemma 5.1.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
$\int_{M}\mathrm{B}(\nabla f,\nabla f)e^{-f}\,\mathrm{d}V=-\frac{1}{2}\int_{M}|\mathrm{D}_{ijk}|^{2}e^{-f}\,\mathrm{d}V.$
Proof.
Let $r$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
$(n-2)\int_{\Omega_{r}}$ $\mathrm{B}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\nabla_{k}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-f}\,\mathrm{d}V-\frac{n-3}{n-2}\int_{\Omega_{r}}\mathrm{C}_{jki}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\nabla_{k}f\cdot|\nabla f|^{-1}\,\mathrm{e}^{-f}\,\mathrm{d}S$
$\quad+\int_{\Omega_{r}}\mathrm{D}_{ikj}(\mathrm{R}_{ik}\nabla_{j}f+\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$\quad-\int_{M}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{D}_{ikj}(\mathrm{R}_{ik}\nabla_{j}f+\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{1}{2}\int_{\Omega_{r}}\mathrm{D}_{ijk}(\mathrm{R}_{ik}\nabla_{j}f-\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{(n-2)}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
6 Proof of Theorem 1.4
First observe that for any $(\alpha,\beta)\in\mathcal{C}$, we can write:
$\mathfrak{B}_{ij}=2\alpha\,\mathrm{B}_{ij}+\Big(\beta-\frac{\alpha}{3}\Big)\,\mathrm{V}_{ij}.$
Assuming $\mathfrak{B}_{ij}=0$ and taking a trace we have:
$g^{ij}\,\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0\implies\Delta\mathrm{R}=0,$
by Lemma 2.1. Then we have:
$\int_{M}\mathfrak{B}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\alpha\int_{M}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{4}\Big(\beta-\frac{\alpha}{3}\Big)\int_{M}\mathrm{R}^{2}\,|\nabla f|^{2}\,\mathrm{e}^{-f}\,\,\mathrm{d}V=0.$
Since $(\alpha,\beta)\in\mathcal{C}$, if $\alpha>0$ and $\beta>\alpha/3$ then both integrals are nonnegative and must each vanish, so $|\mathrm{D}_{ijk}|^{2}\equiv 0$ and $\mathrm{R}^{2}|\nabla f|^{2}\equiv 0$. We conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1. The argument for $\alpha<0$ and $\beta<\alpha/3$ is analogous.
7 The Vanishing U-tensor
The analysis we used for the vanishing $V$-tensor doesn’t completely carry over to $\mathrm{U}=0$, but will allow us to prove Theorem 1.2. Define a weighted volume element of the form
$\mathrm{d}V_{c}(\mathbf{x})=\mathrm{e}^{-cf(\textbf{x})}\,\mathrm{d}V\quad\mathrm{where}\quad c>0.$
Lemma 7.1.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Let $r$ be a regular value of $f$ and define the sub-level set $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. If $c>0$ is a constant and $\Delta\mathrm{R}=0$, then we have the following:
  1. (1.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=0$,
  2. (2.)
    $\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$,
  3. (3.)
    $\int_{\Omega_{r}}f\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=0$.
Proof.
First we have:
$0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S+c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V$
$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V,$
showing (1). Next, we have:
$0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-c\,\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\Big)\mathrm{e}^{-cf}\,\mathrm{d}V.$
Hence, using Lemma 2.2(3),
$\int_{\Omega_{r}}\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
showing (2). And the last identity is:
$\int_{\Omega_{r}}f$ $\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\frac{1}{c}\int_{\Omega_{r}}\langle f\,\nabla\mathrm{R},\nabla(\mathrm{e}^{-cf})\rangle\,\mathrm{d}V$
$=-\frac{1}{c}\Big[\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\mathrm{div}(f\nabla\mathrm{R})\mathrm{e}^{-cf}\,\mathrm{d}V\Big]$
$=-\frac{1}{c}\Big[r\,\mathrm{e}^{-cr}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-cf}\,\mathrm{d}V\Big]$
$=\frac{1}{c}\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\Delta\mathrm{R}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V$
$=0,$
showing (3). ∎
We first state a quick consequence of Bochner’s formula.
Lemma 7.2.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\Delta\mathrm{R}=0$, then:
$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle.$
Proof.
By Bochner’s formula:
$\frac{1}{2}\Delta|\nabla f|^{2}$ $=\langle\nabla(\Delta f),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\mathrm{Rc}(\nabla f,\nabla f)$
$\frac{1}{2}\Delta(f-\mathrm{R})$ $=\langle\nabla(2-\mathrm{R}),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$
$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+\left|\frac{1}{2}g-\mathrm{Rc}\right|^{2}$
$1-\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+1-\mathrm{R}+|\mathrm{Rc}|^{2}$
$\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\mathrm{Rc}|^{2}.$
From here our claim clearly follows. ∎
Lemma 7.3.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Moreover, this can be rewritten as
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\tfrac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
Proof.
One can easily check that on a soliton we have
$\mathrm{U}_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
as well as the soliton identities
  1. (a)
    $\mathrm{Rc}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{Rc}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
  2. (b)
    $\Delta f=2-\mathrm{R}$,
  3. (c)
    $\nabla^{2}f(\nabla f)=\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}$.
The second term in our expression for $\mathrm{U}$ is simplified by looking at the following divergence:
$\nabla_{k}(\mathrm{R}_{ij}$ $\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf})$
$=\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\mathrm{e}^{-cf}+2\,\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}\nabla_{k}f(\nabla_{k}f)\mathrm{e}^{-cf}$
$\quad+\Delta f\,\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-cf}-c|\nabla f|^{2}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf}.$
Now by the divergence theorem we have
$\int_{\Omega_{r}}$ $\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{div}\left(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf}\right)\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
$\quad-\int_{\Omega_{r}}\langle\nabla\mathrm{R},\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\frac{1}{2}\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S+\frac{1}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
So the weighted integral of $\mathrm{U}_{ij}$ becomes
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Now transforming using the Bochner formula from Lemma 7.2:
$\int_{\Omega_{r}}$ $\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}$
$=\frac{1}{4}\int_{\Omega_{r}}\Big(\mathrm{R}^{2}|\nabla f|^{2}-\mathrm{R}|\nabla f|^{2}+\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{d}V_{c}$
$=\frac{1}{4}\int_{\Omega_{r}}\Big(\mathrm{R}^{2}|\nabla f|^{2}-\mathrm{R}|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
The issue is that $\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}$ can change sign over $M$, so
$\int_{M}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}=0$
does not yield obvious rigidity. For comparison, we state the analogous identity for $\mathrm{V}(\nabla f,\nabla f)$ under $\mathrm{d}V_{c}$. The contrast with Lemma 7.3 is instructive: for $c\geq 1$ the integrand for $\mathrm{V}$ acquires a nonnegative term, while the integrand for $\mathrm{U}$ remains sign-indefinite for any $c>0$. This further highlights the subtleties in the $\mathrm{U}=0$ case.
Lemma 7.4.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. If $\Delta\mathrm{R}=0$ and $c>0$, then
$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
and in particular
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}.$
Proof.
We compute the following divergence:
$\nabla_{i}$ $(\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf})$
$=\nabla_{i}\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-cf}+\Delta f\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}$
$\quad+\nabla^{2}f(\nabla f,\nabla\mathrm{R})\mathrm{e}^{-cf}-c|\nabla f|^{2}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}.$
Hence the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
$\int_{\Omega_{r}}$ $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\nabla_{i}(\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{i}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\int_{\Omega_{r}}(\mathrm{R}-2)\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\int_{\Omega_{r}}\Big(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle-\frac{1}{2}|\nabla\mathrm{R}|^{2}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V+c\int_{\Omega_{r}}(f-\mathrm{R})\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S$
$\quad+(c-1)\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\Big(\frac{c-1}{c}+\frac{1}{2}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V.$
Therefore:
$\int_{\Omega_{r}}$ $\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
Taking the limit $r\rightarrow\infty$ gives the stated global identity. ∎
8 Proof of Theorem 1.2
Proof.
Assume that $\mathrm{U}_{ij}=0$. By Lemma 2.1, $\Delta\mathrm{R}=0$. Let $p$ be the point where $f$ attains its minimum value $f_{\min}$. Then $\nabla f(p)=0$, so by (2), $\mathrm{R}(p)=f_{\min}$. Using $\mathrm{U}=0$ at $p$ with $\Delta\mathrm{R}=0$:
$\mathrm{Rc}(p)\big(1-\mathrm{R}(p)\big)+\tfrac{1}{4}\mathrm{R}(p)\big(\mathrm{R}(p)-1\big)g(p)=0,$
which factors as
$\big(1-\mathrm{R}(p)\big)\!\left(\mathrm{Rc}(p)-\tfrac{1}{4}\mathrm{R}(p)\,g(p)\right)=0.$
If $\mathrm{R}(p)\neq 1$ then $\mathrm{Rc}(p)=\tfrac{1}{4}\mathrm{R}(p)g(p)$, which gives
$|\mathrm{Rc}(p)|^{2}=\tfrac{1}{4}\mathrm{R}^{2}(p).$
Since $\nabla f(p)=0$, Lemma 7.2 gives $2|\mathrm{Rc}(p)|^{2}=\mathrm{R}(p)$. Substituting:
$\tfrac{1}{2}\mathrm{R}^{2}(p)=\mathrm{R}(p)\implies\mathrm{R}(p)\big(\mathrm{R}(p)-2\big)=0.$
Hence $\mathrm{R}(p)\in\{0,2\}$. Together with the $\mathrm{R}(p)=1$ case, we have $f_{\min}=\mathrm{R}(p)\in\{0,1,2\}$.
Case $f_{\min}=0$. We have $\mathrm{R}(p)=0$, and the manifold is isometric to the Gaussian soliton by [8].
Case $f_{\min}=2$. Using the soliton equation,
$\mathrm{R}^{2}-2|\mathrm{Rc}|^{2}=\mathrm{R}^{2}-2\left|\tfrac{1}{2}g-\nabla^{2}f\right|^{2}=\mathrm{R}^{2}-2\mathrm{R}+2-2|\nabla^{2}f|^{2}.$
At the minimum $p$ with $f_{\min}=2$, we have $\Delta f(p)=2-\mathrm{R}(p)=0$. Since $p$ is a minimum, $\nabla^{2}f(p)\geq 0$, and $\Delta f(p)=0$ forces $\nabla^{2}f(p)=0$. Thus:
$\mathrm{R}^{2}(p)-2\mathrm{R}(p)+2-2|\nabla^{2}f(p)|^{2}=4-4+2-0=2>0.$
By continuity, there exist $r>2$ and $\epsilon>0$ such that $\mathrm{R}^{2}-2\mathrm{R}+2-2|\nabla^{2}f|^{2}\geq\epsilon>0$ on $\Omega_{r}$. Therefore, under $\mathrm{U}\equiv 0$,
$\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V=0$
implies $|\nabla f|^{2}\equiv 0$ on $\Omega_{r}$, hence $\mathrm{R}=f-|\nabla f|^{2}=2$ on $\Omega_{r}$. By unique continuation for harmonic functions, $\mathrm{R}=2$ on all of $M$, and the soliton is Einstein by [8].
For the converse, if $M$ is Einstein then $\mathrm{B}\equiv 0$ (as $\mathrm{B}$ vanishes on Einstein manifolds [4]) and $\mathrm{V}\equiv 0$ by Theorem 1.1, so $\mathrm{U}=2\mathrm{B}-\tfrac{1}{3}\mathrm{V}\equiv 0$. If $M$ is Gaussian then all curvature vanishes and $\mathrm{U}\equiv 0$.
Case $f_{\min}=1$. At a minimum point $p$ the $\mathrm{U}$-tensor vanishes independently of any global assumption, and $\mathrm{V}$ cannot vanish. Coupled with the sign issue for the $\mathrm{U}$-vanishing integral identity, this case remains open; see Conjecture 1.3. ∎
9 Variational Origins of U and V
The classical Bach tensor arises naturally in a number of ways. One approach is to start with the conformal Einstein equations and then study integrability conditions via divergence operations. A necessary condition to be conformally Einstein is for the Bach tensor $\mathrm{B}_{ij}$ to vanish. The approach taken by R. Bach [2] was to consider the first variation of the quadratic Riemannian functional
$\mathrm{S}_{W}=\int|\mathrm{W}|^{2}\,\,\mathrm{d}V_{g}.$
As seen in A. Besse [3], one has
$\mathrm{B}\,\,\propto\,\,\mathrm{grad}\left(\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}\right),$
so the Bach tensor is the gradient of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. We work in the standard transverse-traceless (TT) gauge for the second variation, imposing for a variation $h$:
$\mathrm{(TT)\,gauge:}\quad\mathrm{Div}_{g}(h)=0\quad\mathrm{and}\quad\mathrm{tr}_{g}(h)=0.$
Recall the Lichnerowicz Laplacian acting on symmetric $2$-tensors:
$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki},$
where $\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h$.
Lemma 9.1.
In dimension $4$, define
$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|\mathrm{W}_{g}|^{2}\,\mathrm{d}V_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}\mathrm{R}_{g}^{2}\,\mathrm{d}V_{g}.$
Then the $L^{2}$-gradient satisfies $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$. In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
Proof.
In dimension four:
$\delta\!\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{B},\delta g\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta\!\int\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{V},\delta g\rangle\,\mathrm{d}V_{g}.$
Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$:
$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.\qed$
Corollary 9.2.
$\mathrm{V}\;=\;-\tfrac{1}{2}\,\mathrm{grad}\!\left(\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}\right),\qquad\mathrm{U}\;=\;-\,\mathrm{grad}\!\left(\int_{M}\big(|\mathrm{W}|^{2}-\tfrac{1}{6}\mathrm{R}^{2}\big)\,\,\mathrm{d}V_{g}\right).$
Proposition 9.3 (Second variation of $\int\mathrm{R}^{2}$ in TT gauge).
Let $(M^{4},g)$ be closed. For a TT variation $g_{t}=g+th+o(t)$, the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}-2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}+2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}+2\langle\mathrm{Rc},h\rangle^{2}+\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,\,\mathrm{d}V.$
Proof.
Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Differentiating once more and using TT gauge:
$\delta^{2}F[h,h]=-2\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
Expanding using the standard TT-gauge variations
$\delta\mathrm{R}$ $=-\langle\mathrm{Rc},h\rangle,$
$\delta(\Delta\mathrm{R})$ $=\Delta(\delta\mathrm{R})-\langle h,\nabla^{2}\mathrm{R}\rangle,$
$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij},$
and integrating by parts yields the result. ∎
Remark 9.4.
If $(M^{4},g)$ is Einstein, the main contributions to the second variation come from the derivative term $\mathrm{R}\,|\nabla h|^{2}$ and the curvature terms involving $\mathrm{R}_{ikjl}\,h_{ij}\,h_{kl}$.
Proposition 9.5 (Second variation at Einstein backgrounds).
On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and $\mathrm{R}=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in TT gauge is
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle\,\mathrm{d}V.$
Equivalently, for each TT eigenmode $\Delta_{L}h_{k}=\mu_{k}h_{k}$,
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\sum_{k}p_{\alpha,\beta}(\mu_{k})\,\|h_{k}\|_{L^{2}}^{2},\quad p_{\alpha,\beta}(\mu)=\alpha\mu^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\mu+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta).$
Proof.
In TT gauge at an Einstein metric:
$\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\qquad\mathrm{V}^{\prime}(h)=\frac{1}{2}\mathrm{R}\,\Delta_{L}h.$
The claimed polynomial follows by expanding
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.\qed$
Corollary 9.6 (Bach line and factorization).
On the Bach line $\beta=\alpha/3$:
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g}.$
Corollary 9.7 (Local rigidity by spectral positivity).
Let $(M^{4},g_{E})$ be Einstein with TT-spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$. If $p_{\alpha,\beta}(\mu)>0$ for every TT-eigenvalue $\mu$, then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling. In particular, there exists a neighborhood of $g_{E}$ in which $\alpha\,\mathrm{U}+\beta\,\mathrm{V}=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
Remark 9.8 (Flat/Gaussian background).
If $\mathrm{R}=0=\mathrm{W}$, then $\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $\mathrm{V}^{\prime}(h)\equiv 0$ on TT. Hence
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V,$
which is strictly positive for $\alpha>0$ on compactly supported variations.